# font=small

### v1

@@ 1 Introduction
Cavity quantum electrodynamics (cQED) has taken center stage with the recent surge in studies showing cavity control of matter-based phenomena, including chemical reactions 1, 2, 3, 4, 5, 6, 7 and energy transfer 8, 9, 10, 11, 12. Early efforts to rationalize such cavity control resorted to minimal quantum representations, such as the Jaynes–Cummings 13 and Tavis–Cummings models 14, but it has since become evident that more elaborate multi-modal representations are necessary 15, 16, 17, 18, 19, 20, 21, 22, 23, 24 to capture collective effects, dark states, disorder, and dissipation 3, 25, 26, 27. In such cases, retaining a full quantum-mechanical description of light and matter becomes prohibitively expensive, prompting the need to take the classical approximation for part of the involved coordinates.
@@
A natural approach to model cQED under a partial classical approximation is by invoking classical electromagnetism, while retaining a quantum description of matter. In principle, this approach unifies established techniques from classical optics, including finite-difference time-domain 28 and finite element methods 29, 30, with those from quantum chemistry. Such unification can be realized through various nonadiabatic dynamics approaches based on quasiclassical 31, 32, 33 or mixed quantum–classical dynamics 34, 35. Within the latter category, a straightforward means to ensure self-consistency of the light–matter interactions is provided by the Ehrenfest theorem, yielding a mean-field (MF) approach to cQED 34, 35.
@@
While MF dynamics is able to broadly capture trends arising from quantum emitters interacting with cavity modes 34, 35, we have previously shown such interactions to lead to a catastrophic dissipation of cavity vacuum fluctuations under the application of the Ehrenfest theorem, yielding qualitatively-incorrect dynamics 36, 37. We have resolved this issue by modifying the Ehrenfest theorem such that vacuum fluctuations are selectively decoupled from the lowest-energy quantum state for all optical transitions of the quantum emitters, upon which near-quantitative accuracy was reached for examples involving two and three-level atoms embedded in a long cavity 36, 37. This modification, however, does not address the short-cavity limit, where mixed quantum–classical dynamics may generally be expected to deliver a poor performance. Very recently, while investigating this performance, we discovered 38 that MF dynamics captures the short-cavity limit remarkably well when the commonly-applied Wigner sampling of the initial classical cavity coordinates 39 is replaced by a focused sampling scheme 40, 41 that (approximately) reinforces zero-point energy (ZPE) at the single-trajectory level. As an additional benefit, we found this approach to reduce the number of trajectories needed to reach converged results from thousands down to as little as one 38.
@@
The abovementioned advances leave us with two distinct approaches reaching high accuracies in the short and long-cavity regimes, respectively, but there remains the need for a single low-cost methodology that reaches desirable accuracy for both short and long cavities. In this Article, we deliver such a methodology by incorporating focused sampling into DC-MF dynamics. While our previous survey of focused sampling required an overshooting of the ZPE in order to quantitatively capture the frequency of recurrent light–matter interactions in cQED 38, we show here that a renormalization of the matter-based transition dipole moments (TDMs) can be applied to the same effect. Our results for DC-MF dynamics show the renormalization of TDMs to be a preferred approach. Altogether, this yields a methodology that optimally balances computational efficiency and accuracy across cavity lengths.
2 Short-Cavity Regime
@@
We will first proceed to briefly summarize MF and DC-MF dynamics, as well as the employed model system, referring to our previous work for details 36, 37. We consider an atom interacting with the field of an optical cavity, as governed by the total mixed quantum–classical Hamiltonian
@@
where $\alpha$ labels the cavity modes, and $\omega_{\alpha}$ denotes the corresponding angular frequency. Further, $Q_{\alpha}$ and $P_{\alpha}$ denote the corresponding position and momentum coordinates, which are associated with the electric and magnetic field components, respectively 42. As before 43, 34, 36, we consider the cavity to have a reduced dimensionality, invoking one light-propagation direction and one orthogonal polarization direction. Consequently, modes organize in a nondegenerate series, with the corresponding angular frequencies given by $\omega_{\alpha}=\pi c_{0}\alpha/L$. Here, $c_{0}$ is the speed of light and $L$ is the cavity length. The atom–field interaction Hamiltonian operator is then given by
$\hat{H}_{\mathrm{AF}}=\sum_{\alpha}\sum_{k<l}\omega_{\alpha}\lambda_{\alpha}\mu_{kl}Q_{\alpha}\ket{k}\bra{l}+\mathrm{H.c.},$ (5)
@@
is the field amplitude of mode $\alpha$ at the atomic location $r_{\mathrm{A}}$ along the light-propagation direction, with $\epsilon_{0}$ as the vacuum permittivity. Here and henceforth, we assume the atom to reside at the center of the cavity, as a result of which this amplitude reduces to $\lambda_{\alpha}=(2/\epsilon_{0}L)^{1/2}$ and $\lambda_{\alpha}=0$ for odd and even $\alpha$, respectively. The cavity coordinates are propagated using the Hamilton equations, with the force terms derived from the total Hamiltonian as
$F_{\alpha}=-\Braket{\psi}{\frac{\partial\hat{H}}{\partial Q_{\alpha}}}{\psi}.$ (7)
Whereas in MF dynamics the classical coordinates are representative of both vacuum fluctuations and thermal fluctuations, DC-MF dynamics separates those fluctuations into two separate coordinates. Accordingly, the optical field Hamiltonian function is expressed as 36
@@
with the thermal and vacuum positions denoted as $q_{\alpha}$ and $Q_{\alpha}$, respectively, and analogously for the momenta $p_{\alpha}$ and $P_{\alpha}$.aaaWe adopted a change in notation compared to Refs. 36 and 37, where the vacuum and thermal coordinates were expressed as upper case characters with and without tilde, respectively. This notation change simplifies the discussion in the present Article. These coordinates are separately evolved through the Hamilton equations, with force terms given by
@@
For every optical transition, the vacuum positions $Q_{\alpha}$ are then decoupled from the lower level of the atom, yielding an atom–field interaction Hamiltonian of the form 37
$\hat{H}^{\mathrm{DC}}_{\mathrm{AF}}=\sum_{\alpha}\sum_{k<l}\omega_{\alpha}\lambda_{\alpha}\mu_{kl}\left(q_{\alpha}+\rho_{l}Q_{\alpha}\right)\ket{k}\bra{l}+\mathrm{H.c.}$ (10)
Here, the term $\rho_{l}\equiv\braket{\psi}{l}\braket{l}{\psi}$ represents the occupancy of atomic level $l$, and is responsible for the decoupling.
@@
As before 43, 34, 36, we consider the zero-temperature limit, such that thermal fluctuations of the optical field are omitted when initializing the atom–cavity system and only the vacuum fluctuations need to be accounted for. Arguably, the most common approach to initialize classical coordinates representative of vacuum fluctuations is by means of Wigner sampling, which is schematically illustrated for a single cavity mode in Fig. 1 (a,c). Accordingly, $Q_{\alpha}$ and $P_{\alpha}$ are stochastically drawn from the normal distribution
$P({Q_{\alpha},P_{\alpha}})=\frac{1}{\pi\hbar}\mathrm{exp}\left(-\frac{P^{2}_{\alpha}}{\hbar\omega_{\alpha}}-\frac{\omega_{\alpha}Q^{2}_{\alpha}}{\hbar}\right).$ (11)
@@
Under such sampling, the ensemble average reproduces the ground-state quantum harmonic oscillator wavefunction, as stipulated by the Wigner quasiprobability distribution 39. While ZPE is thus reproduced at the ensemble level, single trajectories will initialize that undershoot or overshoot this energy.
@@
Figure 2: Transient excited-state occupancy of a two-level atom embedded in a half-wavelength cavity, resulting from MF(W) (red) and DC-MF(W) (blue) dynamics (a), and from MF(f1.06) (orange) and DC-MF(f1.33) (purple) dynamics (b). Results from CISD are also shown (gray).
@@
Recently, we found MF dynamics under Wigner sampling to perform poorly in the short-cavity limit 38. Specifically, we considered a two-level atom embedded in a cavity whose length was taken to amount to half the wavelength associated with the atom’s optical transition, so that the fundamental mode of the cavity has an angular frequency satisfying $\omega_{1}=(\epsilon_{2}-\epsilon_{1})\hbar^{-1}$. Under such resonant conditions, the atom interacts significantly only with this fundamental mode (so long as the ultra or deep-strong coupling regime is avoided). Absent any dissipation, the transient atomic (and optical) occupancies are then expected to exhibit a persistent Rabi oscillation at a frequency $2g$, where we define $g\equiv\mu_{12}\lambda_{1}\sqrt{\omega_{1}/2\hbar}$, with $\lambda_{1}$ as the field amplitude of the fundamental mode. However, instead of such a persistent Rabi oscillation, MF dynamics under Wigner sampling showed a marked decay 38. We proceed by revisiting this simulation, while referring to the applied methodology as MF(W) dynamics.
@@
The result is reproduced in Fig. 2 (a), where a comparison is drawn of the transient atomic excited-state occupancy, $\rho_{2}$, with full-quantum results produced by configuration interaction singles and doubles (CISD), details of which can be found in Ref. 36. The parameters are chosen such that $\omega_{1}=50g$, where it is noted that the results as presented are otherwise invariant to the choice of parameters. An average of a total of 50 000 trajectories was taken in order to ensure convergence. Beyond revisiting these previously-reported results, we also include in Fig. 2 (a) the outcome of DC-MF dynamics where vacuum coordinates are sampled from a Wigner distribution while thermal coordinates are initialized as zero, which we refer to as DC-MF(W) dynamics. Here, an averaging is taken over a total of 10 000 trajectories. Whereas the oscillation decay is seen to be somewhat less severe compared to MF(W) dynamics, the overall performance is still poor, with a conspicuous underestimation of the overall Rabi oscillation frequency.
@@
Previously 38, we were able to attribute the Rabi oscillation decay observed for MF(W) dynamics to a destructive interference arising from a dependence of the oscillation frequency on the initial energy of the cavity mode. This prompted us to replace the Wigner sampling scheme with one where the mode energy is fixed to the value that reproduces the correct frequency. Such a reinforcement of properties at the single-trajectory level is known as focused sampling 40, 41. We resorted to focused sampling primarily in order to resolve unphysical behavior arising from destructive interference under Wigner sampling, but in many instances such interference leaves observables unchanged. It does, however, yield an unnecessary increase in the computational cost of a simulation. As a result, focused sampling techniques are commonly employed for the sake of computational cost reduction 44, 45, 46, with reported improvement factors lying between 10 (Ref. 47) and 25 (Ref. 48).
@@
The focused sampling approach is schematically illustrated for a single cavity mode in Fig. 1 (b,d). Accordingly, initial values of the classical coordinates are drawn from the phase-space circle associated with a given energy through sampling of an angle $\theta_{\alpha}^{(0)}$, as
@@
with $n^{(0)}$ as the initial (classical) occupancy of the cavity mode. Upon fixing $n^{(0)}=1/2$, the initial mode energy equals $\hbar\omega_{\alpha}/2$, meaning that ZPE is enforced exactly.
@@
Figure 3: Transient excited-state occupancy of the two-level atom embedded in a half-wavelength cavity resulting from MF dynamics (a) and DC-MF dynamics (b) as a function of the initial optical occupancy, $n^{(0)}$. Also shown are the corresponding Fourier amplitudes (c and d). Indicated with dashes are the expected Rabi oscillation frequency at $\omega=2g$ (red) and the optical occupancy corresponding to the ZPE at $n^{(0)}=1/2$ (white).
In Fig. 3 (a), we show the transient excited-state occupancy of the two-level atom resulting from MF dynamics under focused sampling as a function of $n^{(0)}$. Such an evaluation was previously conducted by analytically solving for the mixed quantum–classical dynamics under the Ehrenfest theorem 38. For consistency reasons, here we reproduced these previous results numerically within a MF dynamics simulation, while noting that this simulation converges within a single trajectory, yielding results that are virtually indistinguishable from the previous analytical analog 38. The Fourier amplitude of the transient occupancy is presented in Fig. 3 (c),bbbThe Fourier transform was performed over a duration of $gt=200$, after multiplication of $\rho_{2}$ by an exponential with an $1/e$ time of $gt=20$ to avoid clipping and after subtracting a 0.5 baseline, upon which the signal was normalized. showing a dominant component whose energy increases with $n^{(0)}$, alongside a pair of additional components that vanish with increasing $n^{(0)}$. The additional components are representative of anharmonicities arising for the Rabi model under mixed quantum–classical modeling 38.
@@
All components are seen to reflect the aforementioned dependence of the Rabi oscillation frequency on the initial energy of the cavity mode, which ultimately causes destructive interference under Wigner sampling. The dominant component was previously shown 38 to follow the frequency relationship $\omega\approx 2g(n^{(0)}+1/2)^{1/2}$, but its frequency undershoots this equality in the limit $n^{(0)}\rightarrow 0$, in order to reach $\omega=0$ for $n^{(0)}=0$. The latter reflects that mixed quantum–classical dynamics uniquely predicts a complete absence of emission by the atom when the optical field is unoccupied 49. As a consequence, the dominant component crosses the ZPE at an occupancy slightly exceeding $n^{(0)}=1/2$. Previously, we therefore adopted an adjusted value of $n^{(0)}=0.59$ in order to have MF dynamics accurately reproduce the Rabi oscillations 38.
@@
A similar analysis is presented in Fig. 3 (b,d), but for DC-MF dynamics. Different from MF dynamics, this simulation required an averaging over a total of 10 trajectories in order to ensure convergence. While anharmonicities are apparent, similarly to MF dynamics, there are a number of comparative differences seen in the results. First, the dominant component shows an energy increase with $n^{(0)}$ that is much slower compared to what was found for MF dynamics, yielding a ZPE crossing at $n^{(0)}\approx 6$. Second, anharmonicities do not generally vanish with increasing $n^{(0)}$, but instead are organized in a variety of components, some vanishing and some growing in. Combined, however, these components leave a window around $n^{(0)}\approx 1/2$, where Rabi oscillations are reasonably harmonic. Altogether, these results show that having DC-MF dynamics reproduce Rabi oscillations by adjusting $n^{(0)}$ is not a feasible approach, owing to the large $n^{(0)}$ value necessary, and the strong anharmonicity present at this value.

### v3

@@ 1 Introduction
Cavity quantum electrodynamics (cQED) has taken center stage with the recent surge in studies showing cavity control of matter-based phenomena, including chemical reactions hutchisonModifyingChemicalLandscapes2012, thomasGroundStateChemicalReactivity2016, ribeiroPolaritonChemistryControlling2018, hertzogStrongLightMatter2019, hiraiRecentProgress2020, dunkelbergerVibrationCavityPolaritonChemistry2022, mandalTheoreticalAdvances2023 and energy transfer colesPolaritonmediatedEnergyTransfer2014, zhongNonRadiative2016, rozenmanLongRangeTransport2018, georgiouUltralongRange2021, pandyaTuningCoherentPropagation2022. Early efforts to rationalize such cavity control resorted to minimal quantum representations, such as the Jaynes–Cummings jaynesComparisonQuantumSemiclassical1963 and Tavis–Cummings models tavisExactSolutionMolecule1968, but it has since become evident that more elaborate multi-modal representations are necessary hoffmannEffectManyModes2020, wangRoadmapTheoryVibrational2021, tichauerMultiScale2021, balasubrahmaniyamCouplingDecouplingPolaritonic2021, fregoniTheoreticalChallengesPolaritonic2022, ribeiroMultimodePolaritonEffects2022, sanchez-barquillaTheoreticalPerspectiveMolecular2022, mandalMicroscopicTheoryMultimode2023, liMesoscaleMolecularSimulations2024, sunExploringDelocalizationDark2025 to capture collective effects, dark states, disorder, and dissipation ribeiroPolaritonChemistryControlling2018, scholesEmergence2021, engelhardtUnusualDynamical2022, changDarkState2024. In such cases, retaining a full quantum-mechanical description of light and matter becomes prohibitively expensive, prompting the need to take the classical approximation for part of the involved coordinates.
@@
A natural approach to model cQED under a partial classical approximation is by invoking classical electromagnetism, while retaining a quantum description of matter. In principle, this approach unifies established techniques from classical optics, including finite-difference time-domain yeeNumericalSolutionInitial1966 and finite element methods hrennikoffSolutionProblemsElasticity1941, courantVariationalMethodsSolution1943, with those from quantum chemistry. Such unification can be realized through various nonadiabatic dynamics approaches based on quasiclassical liQuasiclassicalModelingCavity2020, SallerBenchmarkingQuasiclassicalMapping2021, heCommutatorMatrix2021 or mixed quantum–classical dynamics hoffmannCapturingVacuumFluctuations2019, hoffmannBenchmarkingSemiclassicalPerturbative2019. Within the latter category, a straightforward means to ensure self-consistency of the light–matter interactions is provided by the Ehrenfest theorem, yielding a mean-field (MF) approach to cQED hoffmannCapturingVacuumFluctuations2019, hoffmannBenchmarkingSemiclassicalPerturbative2019.
@@
While MF dynamics is able to broadly capture trends arising from quantum emitters interacting with cavity modes hoffmannCapturingVacuumFluctuations2019, hoffmannBenchmarkingSemiclassicalPerturbative2019, we have previously shown such interactions to lead to a catastrophic dissipation of cavity vacuum fluctuations under the application of the Ehrenfest theorem, yielding qualitatively-incorrect dynamics hsiehMeanFieldTreatmentVacuum2023, hsiehEhrenfestModelingCavity2023. We have resolved this issue by modifying the Ehrenfest theorem such that vacuum fluctuations are selectively decoupled from the lowest-energy quantum state for all optical transitions of the quantum emitters. Referred to as decoupled mean-field (DC-MF) dynamics, this approach was shown to reach near-quantitative accuracy for examples involving two and three-level atoms embedded in a long cavity hsiehMeanFieldTreatmentVacuum2023, hsiehEhrenfestModelingCavity2023. However, this left the short-cavity limit unaccounted for, where mixed quantum–classical dynamics may generally be expected to deliver a poor performance. Very recently, while investigating this limit, we discovered hsiehMixedQuantumClassicalDynamics2025 that MF dynamics captures the short-cavity limit remarkably well when the commonly-applied Wigner sampling of the initial classical cavity coordinates wignerOn1932 is replaced by a focused sampling scheme stockFlow1999, bonellaSemiclassicalImplementationMapping2003, dukeSimulatingExcitedState2015 that (approximately) enforces zero-point energy (ZPE) at the single-trajectory level. As an additional benefit, we found this approach to reduce the number of trajectories needed to reach converged results from thousands down to as little as one hsiehMixedQuantumClassicalDynamics2025.
@@
The abovementioned advances leave us with two distinct approaches reaching high accuracies in the short and long-cavity regimes, respectively, but there remains the need for a single low-cost methodology that reaches desirable accuracy for both short and long cavities. In this Article, we deliver such a methodology by incorporating focused sampling into DC-MF dynamics. We show such an implementation to necessitate a renormalization of matter-based transition dipole moments (TDMs) in the quantum–classical interaction Hamiltonian in order for it to reproduce the frequency of recurrent light–matter interactions in cQED. The resulting approach presents itself as a methodology that optimally balances computational efficiency and accuracy across cavity lengths.
2 Theoretical Background
@@
We will first proceed to briefly summarize MF and DC-MF dynamics, as well as the employed model system, referring to our previous work for details hsiehMeanFieldTreatmentVacuum2023, hsiehEhrenfestModelingCavity2023. We consider an atom interacting with the field of an optical cavity, as governed by the total mixed quantum–classical Hamiltonian
@@
where $\alpha$ labels the cavity modes, and $\omega_{\alpha}$ denotes the corresponding angular frequency. Further, $Q_{\alpha}$ and $P_{\alpha}$ denote the corresponding position and momentum coordinates, which are associated with the electric and magnetic field components, respectively pellegriniOptimizedEffectivePotential2015. As before flickAtomsMoleculesCavities2017, hoffmannCapturingVacuumFluctuations2019, hsiehMeanFieldTreatmentVacuum2023, we consider the cavity to have a reduced dimensionality, invoking one light-propagation direction and one orthogonal polarization direction. Consequently, modes organize in a nondegenerate series, with the corresponding angular frequencies given by $\omega_{\alpha}=\pi c_{0}\alpha/L$. Here, $c_{0}$ is the speed of light and $L$ is the cavity length. The atom–field interaction Hamiltonian operator is then given by
[…]
@@
is the field amplitude of mode $\alpha$ at the atomic location $r_{\mathrm{A}}$ along the light-propagation direction, with $\epsilon_{0}$ as the vacuum permittivity. Here and henceforth, we assume the atom to reside at the center of the cavity, as a result of which this amplitude reduces to $\lambda_{\alpha}=\pm(2/\epsilon_{0}L)^{1/2}$ and $\lambda_{\alpha}=0$ for odd and even $\alpha$, respectively. The cavity coordinates are propagated using the Hamilton equations, with the force terms derived from the total Hamiltonian as
$F_{\alpha}=-\Braket{\psi|\frac{\partial\hat{H}}{\partial Q_{\alpha}}|\psi}.$ (7)
Whereas in MF dynamics the classical coordinates are representative of both vacuum fluctuations and thermal fluctuations, DC-MF dynamics separates those fluctuations into two separate coordinates. Accordingly, the optical field Hamiltonian function is expressed as hsiehMeanFieldTreatmentVacuum2023
@@
with the thermal and vacuum positions denoted as $q_{\alpha}$ and $Q_{\alpha}$, respectively, and analogously for the momenta $p_{\alpha}$ and $P_{\alpha}$. (We adopted a change in notation compared to Refs. hsiehMeanFieldTreatmentVacuum2023 and hsiehEhrenfestModelingCavity2023, where the vacuum and thermal coordinates were expressed as upper case characters with and without tilde, respectively. This notation change simplifies the discussion in the present Article.) These coordinates are separately evolved through the Hamilton equations, with force terms given by
$f_{\alpha}=-\Braket{\psi|\frac{\partial\hat{H}}{\partial q_{\alpha}}|\psi},\qquad F_{\alpha}=-\Braket{\psi|\frac{\partial\hat{H}}{\partial Q_{\alpha}}|\psi}.$ (9)
For every optical transition, the vacuum positions $Q_{\alpha}$ are then decoupled from the lower level of the atom, yielding an atom–field interaction Hamiltonian of the form hsiehEhrenfestModelingCavity2023
@@
Here, the term $\rho_{l}\equiv\braket{\psi|l}\braket{l|\psi}$ represents the occupancy of atomic level $l$. It is responsible for a gradual decoupling of vacuum contributions to the lower level in each transition. This decoupling becomes absolute once the atom is in the lowest state associated with a given transition.
@@
As before flickAtomsMoleculesCavities2017, hoffmannCapturingVacuumFluctuations2019, hsiehMeanFieldTreatmentVacuum2023, we consider the zero-temperature limit, such that thermal fluctuations of the optical field are omitted when initializing the atom–cavity system and only the vacuum fluctuations need to be accounted for. Arguably, the most common approach to initialize classical coordinates representative of vacuum fluctuations is by means of Wigner sampling, which is schematically illustrated for a single cavity mode in Fig. 1 (a,c). Denoting initial values of classical coordinates with superscript $(0)$ as a short-hand notation for $t=0$, the initial position and momentum coordinates of the mode are stochastically drawn from the normal distribution
$P({Q_{\alpha}^{(0)},P_{\alpha}^{(0)}})=\frac{1}{\pi\hbar}\mathrm{exp}\left(-\frac{{P^{(0)}_{\alpha}}^{2}}{\hbar\omega_{\alpha}}-\frac{\omega_{\alpha}{Q^{(0)}_{\alpha}}^{2}}{\hbar}\right).$ (11)
@@
Under such sampling, the ensemble average reproduces the ground-state quantum harmonic oscillator wavefunction, as stipulated by the Wigner quasiprobability distribution wignerOn1932. While ZPE is thus reproduced at the ensemble level, single trajectories will initialize that undershoot or overshoot this energy.
3 Short-Cavity Regime
@@
Figure 2: Transient excited-state occupancy of a two-level atom embedded in a half-wavelength cavity, resulting from MF(W) (red) and DC-MF(W) (blue) dynamics. Results from CISD are also shown (gray).
@@
Recently, we found MF dynamics under Wigner sampling to perform poorly in the short-cavity limit hsiehMixedQuantumClassicalDynamics2025. Specifically, we considered a two-level atom embedded in a cavity whose length was taken to amount to half the wavelength associated with the atom’s optical transition, so that the fundamental mode of the cavity has an angular frequency satisfying $\omega_{1}=(\epsilon_{2}-\epsilon_{1})\hbar^{-1}$. Under such resonant conditions, the atom interacts significantly only with this fundamental mode (so long as the ultra or deep-strong coupling regime is avoided). Absent any dissipation, the transient atomic (and optical) occupancies are then expected to exhibit a persistent Rabi oscillation at a frequency $2g$, where we define
$g\equiv\mu_{12}\lambda_{1}\sqrt{\omega_{1}/2\hbar},$ (12)
with $\lambda_{1}$ as the field amplitude of the fundamental mode. However, instead of such a persistent Rabi oscillation, MF dynamics under Wigner sampling showed a marked decay hsiehMixedQuantumClassicalDynamics2025. We proceed by revisiting this simulation, while referring to the applied methodology as MF(W) dynamics.
@@
The result is reproduced in Fig. 2. Here, a comparison is drawn of the transient atomic excited-state occupancy, $\rho_{2}$, with full-quantum results produced by configuration interaction singles and doubles (CISD), details of which can be found in Ref. hsiehMeanFieldTreatmentVacuum2023 and the Supporting Information (SI). The parameters are chosen such that $\omega_{1}=50g$, which falls within the regime where the rotating wave approximation is applicable. Since $\omega_{1}=(\epsilon_{2}-\epsilon_{1})\hbar^{-1}$, and with the transient occupancy depicted as a function of $gt$, the result as shown does not depend on the numerical values taken for the parameters. An average of a total of 50 000 trajectories was taken in order to ensure convergence. Beyond revisiting these previously-reported results, we also include in Fig. 2 the outcome of DC-MF dynamics where vacuum coordinates are sampled from a Wigner distribution while thermal coordinates are initialized as zero, which we refer to as DC-MF(W) dynamics. Here, an averaging is taken over a total of 10 000 trajectories. Whereas the oscillation decay is seen to be somewhat less severe compared to MF(W) dynamics, the overall performance is still poor, with a conspicuous underestimation of the overall Rabi oscillation frequency.
Previously hsiehMixedQuantumClassicalDynamics2025, we were able to attribute the Rabi oscillation decay observed for MF(W) dynamics to a destructive interference arising from a dependence of the oscillation frequency on the initial energy of the cavity mode. This prompted us to replace the Wigner sampling scheme with one where the mode energy is fixed to the value that reproduces the expected Rabi frequency. We refer to such a enforcement of properties at the single-trajectory level as focused sampling, drawing from analogous terminology coined in previous works on photo-induced chemical processes bonellaSemiclassicalImplementationMapping2003, dukeSimulatingExcitedState2015. In case of the latter, sampling was focused so as to ensure that each trajectory was initialized on a single electronic surface bonellaSemiclassicalImplementationMapping2003, dukeSimulatingExcitedState2015. Instead, our applied focusing scheme restricts the cavity mode to a given initial energy. This in turn draws inspiration from previous quasiclassical studies, where restricting modes to the ZPE was found to yield improved accuracy schatzOriginCrossSection1983, luClassicalMechanicsIntramolecular1989, ben-nunZeroPointEnergy1996, stockFlow1999.
@@
Notably, while we resorted to focused sampling primarily in order to resolve unphysical behavior arising from destructive interference under Wigner sampling, in many instances such interference leaves observables unchanged. It does, however, yield an unnecessary increase in the computational cost of a simulation. As a result, focused sampling techniques are commonly employed for the sake of computational cost reduction dukeSimulatingExcitedState2015, mandalQuasiDiabatic2018, runesonSpinMapping2019, ananthPathIntegrals2022, with reported improvement factors lying between 10 (Ref. kimQuantumClassical2008) and 25 (Ref. bonellaLinearizedPath202005).
@@
Our employed focused sampling scheme is schematically illustrated for a single cavity mode in Fig. 1 (b,d). Accordingly, initial values of the classical coordinates are drawn from the phase-space circle associated with a given energy through sampling of an angle $\theta_{\alpha}^{(0)}$, as
@@
with $n^{(0)}$ as the initial (classical) occupancy of the cavity mode. Denoting the initial energy contained in this mode as $E_{\alpha}^{(0)}$, this occupancy equals $E^{(0)}_{\alpha}/\hbar\omega_{\alpha}$. Upon fixing $n^{(0)}=1/2$, the initial mode energy thus equals $\hbar\omega_{\alpha}/2$, meaning that ZPE is enforced exactly.
@@
Figure 3: Transient excited-state occupancy of the two-level atom embedded in a half-wavelength cavity resulting from MF dynamics (a) and DC-MF dynamics (b) as a function of the initial optical occupancy, $n^{(0)}$. Also shown are the corresponding Fourier amplitudes (c,d). Indicated with dashes are the expected Rabi oscillation frequency at $\omega=2g$ (red) and the optical occupancy corresponding to the ZPE at $n^{(0)}=1/2$ (white).
With $\omega_{\alpha}\gg g$, the cavity mode browses the full range of angles well before it begins to exchange energy with the atom. As a result, the atom–field dynamics is independent of $\theta_{\alpha}^{(0)}$, and is dependent only on the initial optical occupancy, $n^{(0)}$. In Fig. 3 (a), we characterize this dependence by showing the transient atomic excited-state occupancy, $\rho_{2}$, as a function of $n^{(0)}$, as calculated within MF dynamics. Previously hsiehMixedQuantumClassicalDynamics2025, we reported a similar result, but there the MF dynamics was solved semi-analytically within the rotating wave approximation. Here, we instead conduct a fully-numerical simulation to ensure consistency with the other results presented in this Article (for which a semi-analytical treatment is not feasible), where it should be noted that the MF dynamics result shown in Fig. 3 (a) is indistinguishable from our previous treatment. We can thus resort to the same arguments as made previously hsiehMixedQuantumClassicalDynamics2025 in order to rationalize the observed trends.
@@
Depicted in Fig. 3 (c) is the Fourier amplitude of the transient occupancy $\rho_{2}$. The underlying Fourier transform was performed over a duration of $gt=200$, after multiplication of $\rho_{2}$ by an exponential with an $1/e$ time of $gt=20$ to avoid clipping and after subtracting a 0.5 baseline, upon which the signal was normalized. The Fourier amplitude shows a dominant component whose energy increases with $n^{(0)}$, alongside a pair of additional components that vanish with increasing $n^{(0)}$. All components are seen to reflect the aforementioned dependence of the Rabi oscillation frequency on the initial energy of the cavity mode, which ultimately causes destructive interference under Wigner sampling. As shown previously hsiehMixedQuantumClassicalDynamics2025, the dominant component follows the frequency relationship $\omega\approx 2g(n^{(0)}+1/2)^{1/2}$, but its frequency undershoots this equality in the limit $n^{(0)}\rightarrow 0$, in order to reach $\omega=0$ for $n^{(0)}=0$. The latter reflects that mixed quantum–classical dynamics uniquely predicts a complete absence of emission by the atom when the optical field is unoccupied liMixedQuantumclassicalElectrodynamics2018. (We note that spontaneous emission is achieved within mixed quantum–classical dynamics for any $n^{(0)}>0$, and that this process does not rely on further quantization of the optical mode.) As a consequence, the dominant component crosses the ZPE at an occupancy slightly exceeding $n^{(0)}=1/2$. Previously hsiehMixedQuantumClassicalDynamics2025, we therefore adopted an adjusted value of $n^{(0)}=0.59$. This value can be inferred from Fig. 3 (c) based on the intersection of the $n^{(0)}$-dependent dominant component with the expected Rabi frequency, $\omega=2g$.
@@
A similar analysis is presented in Fig. 3 (b,d), but for DC-MF dynamics. Whereas the MF dynamics results were obtained using a single trajectory, this simulation required an averaging over a total of 10 trajectories in order to ensure convergence. Similarly to MF dynamics, anharmonicities are apparent, but otherwise there are a number of comparative differences seen in the results. First, the dominant component shows an energy increase with $n^{(0)}$ that is much slower compared to what was found for MF dynamics, yielding a ZPE crossing at $n^{(0)}\approx 6$. Second, anharmonicities do not generally vanish with increasing $n^{(0)}$, but instead are organized in a variety of components, some vanishing and some growing in. Combined, however, these components leave a window around $n^{(0)}\approx 1/2$, where Rabi oscillations are reasonably harmonic. Altogether, these results show that having DC-MF dynamics reproduce Rabi oscillations by adjusting $n^{(0)}$ is not a feasible approach, owing to the large $n^{(0)}$ value necessary, and the strong anharmonicity present at this value.
An alternative and trivial means to affect the Rabi oscillation frequency is through a renormalization of the atomic TDM in the light–matter interaction Hamiltonian. Accordingly, we perform the replacement $\mu_{kl}\leftarrow\gamma\mu_{kl}$ in Eqs. 5 and 10, where $\gamma$ denotes a rescaling factor. This allows one to reproduce the expected Rabi frequency through variations of $\gamma$, while focusing the mode occupancy strictly to $n^{(0)}=1/2$. Not only is this value of $n^{(0)}$ physically-informed, as it enforces ZPE exactly for the mode, it also ensures consistency with analogous focusing approaches applied previously in quasiclassical modeling studies schatzOriginCrossSection1983, luClassicalMechanicsIntramolecular1989, ben-nunZeroPointEnergy1996, stockFlow1999, where it has been argued that any trajectory with mode energies more or less than the ZPE is unphysical. Moreover, at this value the anharmonicities seen in Fig. 3 are minimal. The applicable rescaling factor $\gamma$ can be obtained by a similar intersection procedure applied to Fig. 3 (c,d), and is determined as the ratio of the expected Rabi frequency ($\omega=2g$), divided by the frequency where the dominant component crosses $n^{(0)}=1/2$. This yields $\gamma=1.06$ and $\gamma=1.33$ for MF and DC-MF dynamics, respectively. Conceptually, such a rescaling points to an effective underestimation of the dipolar effect under the applied mixed quantum–classical schemes. (Alternatively, one could associate this rescaling factor with the field amplitudes, $\lambda_{\alpha}$. However, the latter is related to the atomic location along the light-propagation direction, which we expect not to be affected by the partial classical approximation.)
Figure 4: Same as Fig. 2, but for results from MF(f1.06) (orange) and DC-MF(f1.33) (purple) dynamics.